Chapter 1 Problems Saved Help Save & Exit Submit Check my work 10 An owner can lease her building for $120,000 per year for three years. The explicit cost of maintaining the building is $40,000, and the implicit cost is $55,000. All revenues are received, and costs borne, at the end of each year. If the interest rate is 5 percent, determine the present value of the stream of 10 points Instructions: Do not round intermediate calculations. Round your final calculation to two decimal places. Skipped a. Accounting profits. Book Print References b. Economic profits. Prev 10 of 16 Next
Answer #1

We need to use the following formula to calculate the present value:

Present value = Future value/(1+r)n

In which, r is the rate of interest (here, r = 5% = 0.05) and n is the number of years (here, n=1).

A. Accounting profit at the end of the year = revenue - explicit costs = $120,000 - $40,000 = $80,000.

So, the present value of the accounting profit = $80,000/(1+0.05)1 = $80,000/1.05 = $76,190.48

B. Economic profit at the end of the year = accounting profit - implicit costs = $80,000 - $55,000 = $25,000.

So, the present value of the economic profit = $25,000/(1+0.05)1 = $25,000/1.05 = $23,809.52.